OpenStudy (peterg2001):

help please In order to receive credit, you must SHOW ALL YOUR WORK that is necessary to answer the question. Use a Proportion to Solve, show proportion and work! Question 1: What is 25% of 56? Question 2: 16 is what % of 64 Question 3: 12 is 30% of what number

OpenStudy (peterg2001):

@Directrix

OpenStudy (peterg2001):

@hhelpplzzzz @justuu

OpenStudy (peterg2001):

@AlexandervonHumboldt2

OpenStudy (alexandervonhumboldt2):

1. 56*25/100=14

OpenStudy (alexandervonhumboldt2):

2. 64*x/100=16 25%

OpenStudy (alexandervonhumboldt2):

3. x*30/100=12 40

Directrix (directrix):

@peterg2001 A proportion is two equal ratios. I don't see any in this thread. Just saying.

OpenStudy (peterg2001):

tooooo late